Tom who is traveling at a speed of 67 km/hr sees Jerry 178 m away on the road. What is the minimum constant acceleration must Tom have to avoid Jerry assuming he doesn't move.

1
Expert's answer
2022-02-21T12:11:01-0500

Answer

Using third equation of motion

"v^2=u^2+2as"

Putting all values

And acceleration

"a=\frac{u^2-v^2}{2s}\\=\frac{(18.61)^2-0^2}{2\times178}\\=0.97m/s^2"